OpenStudy (anonymous):

x+10 3x-18 A.24 B.14 C.18 D.16

OpenStudy (icalibear):

do you mean x + 10 = 3x - 18? That would make sense but that doesn't

OpenStudy (anonymous):

oops,yes i forgot the equal sign

OpenStudy (anonymous):

Ok so: x+10=3x-18 +18 +18 x+28=3x -x -x 28=2x divide each side by 2 14=x B

OpenStudy (anonymous):

thats not one of the answer choices. i'm confused about this. my teacher isn't in until Friday since she sick

OpenStudy (icalibear):

You could just use process of elimination

OpenStudy (icalibear):

try out each term and see which one fits. there are only four anyway :)

OpenStudy (anonymous):

What do you mean it's not one of your answers???

OpenStudy (anonymous):

That's the right answer.... and it says above in your choices B. 14

OpenStudy (anonymous):

i miss read that. thank you for point it out! i have dyslexia

OpenStudy (anonymous):

Oh no prob! :-)